Table 15-4
A monopolist faces the following demand curve:

Price

Quantity

$30

0

$25

2.5

$20

5

$15

7.5

$10

10

$5

12.5

$0

15

Refer to Table 15-4. In order to maximize total revenues, the monopolist should produce

a.

12.5 units.

b.

7.5 units.

c.

10 units.

d.

5 units.
Answer #1

ANswer

Option b

7.5 units

================

total revenue =P*Q

TR(0)=0*30=0, TR(2.5)=2.5*25=62.5 and so on

Price Quantity TR
30 0 0
25 2.5 62.5
20 5 100
15 7.5 112.5
10 10 100
5 12.5 62.5
0 15 0

The total revenue is maximum when the firm produces 7.5 units of output

Q=7.5 units